OpenStudy (rosales0228):

GIVING MEDAL!!!! Find the solution for this system of equations. Separate the x- and y- values with a comma. x+6y=7 x-y=-7

OpenStudy (rosales0228):

@mia99

OpenStudy (mia99):

$\begin{bmatrix}x+6y=7 \\x-y=-7\end{bmatrix}$

OpenStudy (mia99):

x+6y=7

OpenStudy (mia99):

x+6y-6y=7-6y x=7-6y

OpenStudy (mia99):

[ (7−6y)−y=−7 ]

OpenStudy (mia99):

[ (7−6y)−y=−7 ] 7-7y=-7 7-7y-7=-7-7 -7y=-14

OpenStudy (mia99):

\[\frac{-7y}{-7}=\frac{-14}{-7}\]

OpenStudy (mia99):

y=2

OpenStudy (mia99):

\[\mathrm{For\:}x=7-6y\]

OpenStudy (mia99):

\[\mathrm{Subsititute\:}y=2\]

OpenStudy (mia99):

\[x=7-6\left(2\right)\quad \Rightarrow \quad x=-5\]

OpenStudy (mia99):

\[The\:solutions\:\to\:the\:system\:of\:equationts\:are:\] \[y=2,\:x=-5\]

OpenStudy (mia99):

@Rosales0228